LHC - latest news

LHC schedule can be found here

Slides from our morning meetings here

Luminosity plots here


Final year of Run 2 - another production year. Things are going well despite there still being something at 16L2 - see luminosity plots for the latest. The problem at 16L2 has become known as the Gruffalo - still not fully understood.


An ultimately successful production year somewhat hampered by an accidental air ingress into the beam vacuum in sector 12 (location 16L2) during the preceeding Christmas technical stop.


Progress as of Monday 27th June 2016 - luminosity plot

Sunday 26th June 2016: design luminosity of 1e34 cm-2s-1 reached.

commissioning progress

Tuesday 5th April 2016

Progress has been good. Squeeze to 40 cm performed, optics measured and corrected. Experimental dipoles (LHCb and ALICE) on, separation and crossing-angle bumps on at injection and in the ramp. Collimators set-up at 450 GeV. System commissioning progressing well.

Saturday 26th March 2016

First ramp to 6.5 TeV and into the squeeze.

Friday 25th March 2016

First beam injected. RF capture rapidly done and circulating beams established.



3rd June 2015

First Stable Beams (3 on 3 bunches). Total beam commissioning time 8 weeks 3 days - still a lot of details to sort out but looking good so far at 6.5 TeV.

21st May 2015

First "test" collisions at 6.5 TeV

5th April 2015

First injection Sunday morning, beams around both directions without too many problems. Followed by capture, RF and instrumentation set-up - both beams ciculating in reasonably good condition by the end of the day.

1st April 2015

The high voltage tests yesterday looked fine. The main dipole magnet circuit of sector 34 now needs to be requalified for operation at 6.5 TeV - this should take around 2 days. Another couple of days are then required for final interlock tests with the machine fully available before we can go for beam.

31st March 2015

The attempt to burn away the short yesterday afternoon appears to have gone well. To be confirmed with high voltage tests today.

28th March 2015

Sector 45 is finally trained to the required level for 6.5 TeV operation after around 50 training quenches. Machine checkout continues to stress test all systems in preparation for beam.

However, we still have a short to ground in sector 34, almost certainly due to a small piece of metallic debris in what is now a well established location. An attempt will be made early next week to vaporise the offending object by injectly a current pulse locally.



24th March 2015

The best-laid schemes o' mice an' men
Gang aft agley

The machine checkout is progressing well. Sector 45's main dipole circuit is slowly getting there and has been trained to a current of 10,875 A (target 11,080 A).

However, during the ramp-down of sector 34 following a training quench on Saturday morning, an earth fault developed in the main dipole circuit. All protection systems functioned properly and no harm was done. The fault has been localized. Further measurements are ungoing before deciding on the strategy to be adopted but there will be certainly be some delay to programme.

21st March 2015

Machine checkout of all systems underway (magnet circuits plus collimators, RF, interlock system, controls, instrumentation etc.).

Training still ongoing in sector 34 (very nearly there) and sector 45 (taking longer than expected - see plot below). We will not put beam in until all sectors are fully trained to 6.5 TeV. This may well push first beam by a couple of days.

quench training


17th March 2015

6 sectors fully trained, last 2 (S34, S45) in progress. S45 is putting up a fight.

12th March 2015

Start of beam commissioning now scheduled for Wednesday 25th March (at the earliest!).

7th March 2015

The sector test is ongoing. Beam has been injected in both directions. The clockwise beam has made it through ALICE and sector 23, the anti-clockwise beam through LHCb and sectors 78 and 67 to the beam dump in point 6. Usual mix of problems to solve but so far things look reasonable.

Powering tests will resume on Monday 9th March. Five sectors fully trained to 6.5 TeV, sector 78 nearly there, sectors 34 and 45 should start soon.


27th February 2015

5 sectors now at 6.5 TeV. Sector 23 should be finished with one more quench, sector 78 has just started training, and preparatory tests are ongoing in sectors 34 and 45.

On the weekend of 7th/8th March we hope to perform a sector test with low intensity bunches and bring the clockwise beam into sector 23, and the anti-clockwise beam into sectors 78 and 67.



15th February 2015

Steady progress with powering tests and quench training of main dipoles (see below). Three sectors are now fully trained to 6.5 TeV. One is in progress, the rest will come soon. Updated schedule has first beam on 23rd March 2015.

RB quenches



5th December 2014

The machine is nearly fully cold. The powering tests of the magnet circuits are in full swing. The big issue here is the need to train the superconducting dipoles to the 6.5 TeV current level. Training involves ramping up the dipoles in a given sector until a quench (transition to the resistive state) of a single dipole is detected. Protection systems swing into action, energy is extracted from the circuit, the current ramps down, and we go around the loop again. Next time round the quenched magnet now holds the current (i.e. it has trained), we go to a higher current and another one of the 154 dipoles in the circuit quenches... 11080 kA is the target current (6.5 TeV with a bit of margin).

Progress is nicely illustrated here. Or below (from 11th December) if you can't access the site.



4th July 2014: The splice consolidation campaign has finished after a remarkable effort from all involved. Sector 67 is at 20 K awaiting electrical continuity tests. Sector 81 is in cool-down (average temperture 210 K). Pressure tests, power converter tests, access system tests etc. ongoing elsewhere in the ring.

Linac2, Booster and PS have restarted with beam.


The first 6 weeks of 2013 were devoted to a proton-lead run. Four weeks of magnet tests led into long shutdown 1 (LS1). During LS1 there will be consolidation and upgrades of all major systems. One principal activity is the opening of all magnet interconnects, the measurement of the resistance of all splices in the cables that carry the current between the main dipole and quadrupole magnets, and the repair of any non-conformities.. There are around 10,000 splices in all.

Recommissioning with beam is foreseen for the start of 2015.

During operations with beam, there is a meeting every morning (8:30 during the week, 9:00 Saturday and Sunday). A record of these meetings can be found below.